TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.BorisBellesia on 2006-12-13 - 20:28

60A converter tests - Wed. 13th Dec.

JPEG P1000122.JPG:
P1000122.JPG

Advancement of today:

  • Pre-connection of the cable of the current leads: on C28-C32_L8
  • All 60A cables disconnected PC side:
  • Problem on the short circuit insulation on the 60A cables in R7 (see pic)

Tomorrow activities:

  • TS-EL, resolution of the problem on the short circuit insulation in R7, no problems for coactivities.
  • if it is possible: El_QA.doc 120A circuits in RR77 and UA83, no problems for coactivities.

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-13 - 17:25

RAT Preparation Cool Down sec78 - Dec 13

PDF RAT_20061213.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x13x17x22/RAT_20061213.pdf

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 12.12.06 Short-to-ground in QBBI.A21R7 (Lyra side) (D.Bozzini, F.Seyvet)
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum
  • 12.12.06 One Pressure Gauge in DFBs cannot be read (F.Millet)
Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 12.12.06 Repositioning of 1 DC cable on DFBAO and of 2 DV cables on DFBAN (all 13kA) (J.C.Guillaume)
  • 08.12.06 DFBXG 100V test on current leads (S.Feher)
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

MircoCoccoli - 13 Dec 2006 - 17:22

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-12-01 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login